OpenStudy (shelby1290):

Determine the value(s) of k such that each trinomial is a perfect square. a) x^2 + 4x + k b) 4x^2 - 12x + k

OpenStudy (misty1212):

two steps only

OpenStudy (misty1212):

what is half of 4?

OpenStudy (shelby1290):

@misty1212 half of 4 is 2

OpenStudy (shelby1290):

what's next?

OpenStudy (anonymous):

\[b^2-4ac=0\]

OpenStudy (anonymous):

\[4^2-4(1)(k)=0\]

OpenStudy (anonymous):

k=?

OpenStudy (shelby1290):

12?

OpenStudy (anonymous):

Try again

OpenStudy (shelby1290):

4^2-4(1)(k) 16-4k=0 divide both sides by -4 k=-4

OpenStudy (shelby1290):

is that right?

OpenStudy (anonymous):

\[16-4k=0 \\ \\ 16=4k \\ \\ k=?\]

OpenStudy (shelby1290):

k = 4 oops

OpenStudy (shelby1290):

thank you

OpenStudy (shelby1290):

now how about part b

OpenStudy (anonymous):

Do the same for the next one \[4x^2 - 12x + k=0 \\ Using \\ ax^2+bx+c=0 \\ \\ b^2-4ac=0\]

OpenStudy (anonymous):

\[(-12)^2-4(4)(k)=0\]

OpenStudy (shelby1290):

-144-16k=0 -144=16k divide both sides by 16 k=-9 ?

OpenStudy (anonymous):

\[(-12)^2= (-12)(-12)=12 \times 12=144\]

OpenStudy (anonymous):

k=9

OpenStudy (shelby1290):

oh i see it's because i didn't put (-12)^2 in brackets

OpenStudy (shelby1290):

thanks for the help though!